OpenStudy (anonymous):

In set-builder notation, how do you write the solutions of the inequality? 3x 10 *>=* 4 A) {x|x*>=*-2} B) {x|x*<=* 2} C) {x|x*<=*-2} D) {x|x*>=* 2}

OpenStudy (dtan5457):

3x 10 *>=* 4 can you clarify that?

OpenStudy (anonymous):

*>=* is greater than or equal to

OpenStudy (dtan5457):

3x 10 what does that mean?

OpenStudy (anonymous):

3 with the letter x , not multiply

OpenStudy (dtan5457):

what's the operation with 10 then?

OpenStudy (anonymous):

oh whoops its 3x + 10 *>=* 4

OpenStudy (dtan5457):

\[3x+10\ge4 \]

OpenStudy (dtan5457):

correct?

OpenStudy (anonymous):

yea

OpenStudy (dtan5457):

alright so you just solve for x here. 3x+10>=4 subtract the 10 3x>=-6 divide by and you get..?

OpenStudy (anonymous):

1x*>=*=-3

OpenStudy (dtan5457):

-6/3=?

OpenStudy (anonymous):

-2

OpenStudy (dtan5457):

so it should be x>=-2

OpenStudy (anonymous):

so A ?

OpenStudy (anonymous):

@dtan5457

OpenStudy (dtan5457):

yes